         Usage of hvdashln.sty

Nakashima

Nov. 16, 1992

Definitions of horizontal and vertical dashed lines for array and tabular environment. Instead of partial horizontal dashed line by the \cline{2-3} by Isozaki, \hdashline fills the full width. The proportion of the dashed line is de-

fined by the two parameters: \hdashlinewidth=2pt and \hdashlinegap=2pt,

where the former defines the width of the dash and the latter sets their gap. The default values of these are 4pt while 2pt is employed in the examples below.

[ \begin{matrix}A & B & C \\ D & E & F \\ G & H & I \end{matrix} ]

a b c
a b c
a b c

a b c
a b c
a b c